OpenStudy (anonymous):

In simplest radical form, what is the mean proportional between 4 and 12? (1) sqrt{48} (2) 4sqrt{3} (3)3sqrt{4} (4) 2sqrt{8}

OpenStudy (anonymous):

let it be x \[\frac{ x }{ 4 }=\frac{ 12 }{x }\] \[x ^{2}=4*12=48,calculate x\]